OpenStudy (anonymous):

solve the following equation for x. if there is no solution, state so. 1/(5x-1)-1/(2x^2+1)=0

OpenStudy (campbell_st):

well start by rewriting the equation \[\frac{1}{5x -1} =\frac{ 1}{2x^2 + 1}\] next cross multiply

OpenStudy (campbell_st):

you will need to get all the terms on the left hand side and you should have a quadratic equation that can be factorised. I forgot to say you should check the denominators and see which values x cannot take.

OpenStudy (anonymous):

its not an equal sign, its a minus sign in the middle of the equation